23. (2 marks) We have a discrete probability distribution where Pr(X = 1) = 0.2,
Pr(X = 2) = 0.3 and Pr(X = 3) = 0.5. Determine μ.
Answer #1

Solution :

Mean = = X * P(X)

= 1 * 0.2 + 2 * 0.3 + 3 * 0.5

= ( 0.2+ 0.6 +1.5 )

= 2.3